# Kinematics of Circumgalactic OĀ VIĀ Gas and Disk Rotation of Star-forming Galaxies

### v1

@@ Abstract
Quasar sightline observations reveal that low-ionization-state gas corotates with the galaxy disk and often at sub-centrifugal velocities, suggesting that the gas is spiraling towards the galaxy disk. However, while observations ubiquitously detect OĀ VIĀ absorption around low-redshift, Ā star-forming galaxies, the relationship between OĀ VIĀ and the galaxy disk, especially the kinematics, is not well-established. This work focuses on the OĀ VIĀ kinematics and its comparison with that of the low ions and galactic disk rotation. We present observations of 18 pairs of quasars and star-forming galaxies. All quasar sightlines intersect the circumgalactic medium (CGM) within 45āĀ from the galaxy major axes. We show that while individual OĀ VIĀ velocity components do not correlate with disk rotation, the bulk of OĀ VIĀ gas in individual sightlines rarely counter-rotates. We then match OĀ VIĀ velocity components with those of low ions by minimizing the difference of their velocity centroids. The OĀ VIĀ velocity components with successful low-ion matches are typically found at small sightline impact parameters and are more likely to corotate with the disk. We suggest that the low-ion-matched OĀ VIĀ velocity components trace the gas co-spatial with the low ions near the extended disk plane in the inner CGM, whereas those without low-ion matches represent the gas at large 3D radii. While the gas at large radii is theoretically expected to kinematically correlate with the disk angular momentum, this correlation is expected to be weaker due to the higher turbulent to mean rotation velocity ration at large radii, consistent with our results.
@@ 1 Introduction
The ubiquitous detection of OĀ VI 1031, 1037Ā absorption around low-redshift, Ā star-forming galaxies in contrast to the low detection rate around quiescent galaxies have received significant attention (Tumlinson etĀ al., 2011; Tchernyshyov etĀ al., 2023). Dwarf galaxies also show the same contrast (Dutta etĀ al., 2024), even though the OĀ VIĀ detection rate and column density of star-forming dwarf galaxies are lower than that of Ā galaxies (Johnson etĀ al., 2015, 2017; Tchernyshyov etĀ al., 2022; Mishra etĀ al., 2024). The origin of this āOĀ VIĀ dichotomyā remains controversial. AGN feedback possibly leads to a low CGM mass fraction, resulting in the low OĀ VIĀ detection rate of quiescent galaxies (Suresh etĀ al., 2017; Nelson etĀ al., 2018; Davies etĀ al., 2020). Some work argued that the quiescent galaxies have higher average halo masses than the star-forming galaxies in the observed galaxy samples, and hence, the higher virial temperatures promote oxygen to higher ionization states. In particular, the predicted CGM properties change significantly for galaxies with halo masses around Ā due to virialization (Oppenheimer etĀ al., 2016; Mathews & Prochaska, 2017; Stern etĀ al., 2018; Sultan etĀ al., 2025). However, recent work with a mass-controlled galaxy sample still found the āOĀ VIĀ dichotomyā. This implies a potential connection between OĀ VIĀ and ongoing star formation, for which the transformation of CGM happens concurrently with the quenching of galaxies (Tchernyshyov etĀ al., 2023).
@@
In addition to the spatial distribution, kinematics provides crucial information for understanding the origin and properties of the CGM traced by OĀ VIĀ and low-ionization-state ions. While OĀ VIĀ absorption has a larger velocity spread compared to that of the low ions, they typically span similar velocity ranges (Werk etĀ al., 2016; Nielsen etĀ al., 2017). However, while the absorption of low ions becomes broader and stronger towards the galaxy minor axes, a result attributed to gas kinematically disturbed by outflows (Bordoloi etĀ al., 2011; Martin etĀ al., 2019; Schroetter etĀ al., 2019), the OĀ VIĀ velocity spread does not vary with the sightline azimuthal angle and galaxy disk inclination (Nielsen etĀ al., 2017). Furthermore, for sightlines near the galaxy major axes, the cool gas (traced by MgĀ II) corotates with the galaxy disk and often at sub-centrifugal velocities (Steidel etĀ al., 2002; Kacprzak etĀ al., 2010, 2011; Ho etĀ al., 2017; Zabl etĀ al., 2019), suggesting that the gas is infalling (Ho & Martin, 2020). On the other hand, the lack of correlation between OĀ VIĀ kinematics and disk rotation potentially implies that OĀ VIĀ alone does not serve as a useful tracer for galaxy gas accretion (Kacprzak etĀ al., 2019, 2025).
On the one hand, with the spatially extended distribution of OĀ VIĀ compared to low ions, it is perhaps unsurprising that the OĀ VIĀ kinematics do not seem to correlate with disk rotation as the low-ionization-state counterparts. In particular, because the CGM angular momentum remains roughly constant with radius, the outer CGM is expected to have a lower mean rotation velocity, resulting in a weaker correlation with the disk rotation. But on the other hand, in contrast to the large number of literatures that observationally confirmed the corotation between MgĀ IIĀ gas and the galaxy disk, only very few work has performed this kinematics comparison with OĀ VIĀ (Kacprzak etĀ al., 2019; Nateghi etĀ al., 2024; Kacprzak etĀ al., 2025).
In this paper, we build a sample of star-forming galaxies with quasar sightlines near the galaxy major axes and exclude galaxies with face-on disks. With quasar spectroscopy from the Cosmic Origin Spectrograph (COS) on the Hubble Space Telescope (HST) and galaxy rotation curve measurements, we analyze and compare the kinematics of the CGM traced by OĀ VIĀ and low ions (e.g., SiĀ II, SiĀ III) with the galaxy disk rotation. We describe the sample selection, data reduction, and measurements in SectionĀ 2. In SectionĀ 3, we present the results of the kinematics comparison among OĀ VIĀ gas, low-ionization-state gas, and disk rotation. We interpret our results and discuss their implications in SectionĀ 4. Finally, SectionĀ 5 summarizes the conclusions. Throughout the paper, we adopt the cosmology from Planck Collaboration etĀ al. (2016), with , , , and .
@@ 2 Data and Observations
We present the analyses of 18 galaxyāquasar pairs to study the kinematics of OĀ VIĀ and low-ionization-state absorbers and their relationship with galaxy disk rotation. We conducted quasar spectroscopy with HST/COS, including new spectra from our HST program and archival observations. We also acquired galaxy spectra from multiple ground-based telescopes. We describe our sample selection in SectionĀ 2.1. SectionĀ 2.2 discusses the galaxy spectroscopic observations and rotation curve measurements, and SectionĀ 2.3 describes the galaxy properties. In SectionĀ 2.4, we describe the HST/COS observations and measurements of absorption-line systems.
@@ 2.1 Sample of GalaxyāQuasar Pairs
Our study focuses star-forming galaxies with quasar sightlines near the major axes of the galaxies. The galaxyāquasar pairs were selected from published literature with known galaxy disk orientation; the disk geometry was obtained either from published work or from the Sloan Digital Sky Survey (SDSS; Blanton etĀ al., 2017). We restricted our sample to galaxies with inclined disks of inclination angles of ā, and each sightline lies within 45āĀ from the galaxy major axis, i.e., with azimuthal angle ā.
The compiled sample consists of 18 galaxyāquasar pairs with a median galaxy redshift of . TablesĀ 1 and 2 specify the coordinates and properties of galaxies, respectively, and TableĀ 3 lists the coordinates of the paired quasars. FigureĀ 1 shows the location of the quasar sightlines relative to the projected major and minor axes of the galaxies on the sky. The impact parameter of the quasar sightlines span between 20 and 280 kpc. The majority of the sightlines intersect the galaxies within 100 kpc, which corresponds to less than half of the virial radius () as indicated by the colors of the markers (see SectionĀ 2.3).
@@ 2.2.1 MMT Blue and Red Channel Observations
We conducted new galaxy spectroscopic observations using the MMT Blue and Red Channel Spectrographs. On February 28, 2020, we observed the galaxies with the Blue Channel and configured it with the 832 lines mm-1Ā grating blazed at 7790 Ć . On January 9, 2021, we used the Red Channel with the 1200 lines mm-1Ā grating blazed at 7700 Ć . For both observations, we applied the blue blocking filter R-63 to block the higher order light. We used the default detector binning, which provided a pixel scale of 03 for both channels and a wavelength dispersion of 0.72 Ć Ā pixel-1Ā and 0.80 Ć Ā pixel-1Ā for the Blue and Red Channel, respectively. We observed the target galaxies by aligning the 10-wide long-slit with the galaxy major axes. For both channels, the 10-wide slitlet provided a spectral resolution of 75ā85 km s-1Ā full width at half maximum intensity (FWHM).
@@
From the two-dimensional galaxy spectra, we measured the galaxy rotation curves from the emission lines (SectionĀ 2.2.4). We also extracted an one-dimensional spectrum for each galaxy and determined the galaxy systemic redshift from the redshifted wavelengths of the , [NĀ II] 6548, 6583, and [SĀ II] 6716, 6731Ā emission lines.
@@ 2.2.2 Keck Observations
The Keck/LRIS observations were described in detail in Ho etĀ al. (2017) and Martin etĀ al. (2019). The LRIS double spectrograph was configured using the D500 dichroic, 1200 lines mm-1Ā grism blazed at 3400Ć Ā (LRISb) and the 900 lines mm-1Ā grism blazed at 5500Ć Ā (LRISr). The galaxy emission lines for rotation curve measurements were captured by LRISr. The detectors were binned by two in both spatial and dispersion directions, which produced a spatial scale of and a dispersion of 1.06 Ć Ā pixel-1. The 10-wide long-slit was either aligned with the position angle of the major axis of the target galaxies, or masks of 10-wide slitlets were designed to include nearby galaxies in the field. The 10-wide slitlets provided an average resolution of 75ā105 km s-1Ā FWHM. The data reduction procedures for the LRISr spectra were explained in Ho etĀ al. (2017) and were similar to those for reducing the MMT spectra described in SectionĀ 2.2.1.
The Keck/ESI observations were described in detail in Kacprzak etĀ al. (2019). In brief, the galaxies were observed using the 10-wide, 20ā²ā²-long ESI slit, and the slitlet was aligned with the major axes of individual galaxies. The detector was binned by two in both the spatial and spectral directions. This produced a dispersion of 22 km s-1Ā pixel-1 and pixel scales of 027 to 034 over the echelle orders that covered the galaxy optical emission lines. The resulting spectral resolution from the 10-wide slitlet was 90 km s-1Ā FWHM. Following the data reduction procedures as outlined in Kacprzak etĀ al. (2019), we re-reduced the ESI spectra and re-examined the galaxy emission lines. The new two-dimensional spectra were adopted to measure the galaxy rotation curves (SectionĀ 2.2.4).
@@ 2.2.3 APO DIS Observations
For galaxies J142459+38211 and J155505+362848, Ho etĀ al. (2017) and Martin etĀ al. (2019) conducted long-slit spectroscopy with APO/DIS to measure the galaxy rotation curves from the Ā emission line. The 15-wide long-slit was aligned with the major axis of each galaxy. The blue and red channels of DIS were configured with the 1200 lines mm-1Ā blazed at 4400Ć Ā and 7300Ć , respectively. The Ā emission was detected by the red spectra, which had a spatial scale and dispersion of 040 and 0.58 Ć Ā pixel-1, respectively. The DIS spectra were reduced as discussed in Ho etĀ al. (2017) and followed the same procedures as we reduced the MMT spectra. The spectral resolution was measured using the arclampsā lines as 50 km s-1Ā FWHM.
@@ 2.2.4 Galaxy Rotation Curves
For each galaxy, we project the models onto the slit, convolve these projected models with a Gaussian model for the seeing disk, and compare them to the PV maps measured from the galaxy emission lines. We minimize the fit residual to estimate Ā and for individual galaxies. In the middle column of FigureĀ 2, the seeing-convolved model for each galaxy before and after the projection are shown using the black dashed and cyan solid curves, respectively. TableĀ 2 lists the adopted Ā and for individual galaxies.

### v2

# Kinematics of Circumgalactic OĀ VI Gas and Disk Rotation of Star-forming Galaxies

@@ Abstract
Quasar sightline observations reveal that low-ionization-state gas corotates with the galaxy disk and often at sub-centrifugal velocities, suggesting that the gas is spiraling towards the galaxy disk. However, while observations ubiquitously detect OĀ VI absorption around low-redshift, star-forming galaxies, the relationship between OĀ VI and the galaxy disk, especially the kinematics, is not well-established. This work focuses on the OĀ VI kinematics and its comparison with that of the low ions and galactic disk rotation. We present observations of 18 pairs of quasars and star-forming galaxies. All quasar sightlines intersect the circumgalactic medium (CGM) within 45ā from the galaxy major axes. We show that while individual OĀ VI velocity components do not correlate with disk rotation, the bulk of OĀ VI gas in individual sightlines rarely counter-rotates. We then match OĀ VI velocity components with those of low ions by minimizing the difference of their velocity centroids. The OĀ VI velocity components with successful low-ion matches are typically found at small sightline impact parameters and are more likely to corotate with the disk. We suggest that the low-ion-matched OĀ VI velocity components trace the gas co-spatial with the low ions near the extended disk plane in the inner CGM, whereas those without low-ion matches represent the gas at large 3D radii. While the gas at large radii is theoretically expected to kinematically correlate with the disk angular momentum, this correlation is expected to be weaker due to the higher turbulent to mean rotation velocity ration at large radii, consistent with our results.
@@ 1 Introduction
The ubiquitous detection of OĀ VI 1031, 1037 absorption around low-redshift, star-forming galaxies in contrast to the low detection rate around quiescent galaxies have received significant attention (Tumlinson etĀ al., 2011; Tchernyshyov etĀ al., 2023). Dwarf galaxies also show the same contrast (Dutta etĀ al., 2024), even though the OĀ VI detection rate and column density of star-forming dwarf galaxies are lower than that of galaxies (Johnson etĀ al., 2015, 2017; Tchernyshyov etĀ al., 2022; Mishra etĀ al., 2024). The origin of this āOĀ VI dichotomyā remains controversial. AGN feedback possibly leads to a low CGM mass fraction, resulting in the low OĀ VI detection rate of quiescent galaxies (Suresh etĀ al., 2017; Nelson etĀ al., 2018; Davies etĀ al., 2020). Some work argued that the quiescent galaxies have higher average halo masses than the star-forming galaxies in the observed galaxy samples, and hence, the higher virial temperatures promote oxygen to higher ionization states. In particular, the predicted CGM properties change significantly for galaxies with halo masses around due to virialization (Oppenheimer etĀ al., 2016; Mathews & Prochaska, 2017; Stern etĀ al., 2018; Sultan etĀ al., 2025). However, recent work with a mass-controlled galaxy sample still found the āOĀ VI dichotomyā. This implies a potential connection between OĀ VI and ongoing star formation, for which the transformation of CGM happens concurrently with the quenching of galaxies (Tchernyshyov etĀ al., 2023).
@@
In addition to the spatial distribution, kinematics provides crucial information for understanding the origin and properties of the CGM traced by OĀ VI and low-ionization-state ions. While OĀ VI absorption has a larger velocity spread compared to that of the low ions, they typically span similar velocity ranges (Werk etĀ al., 2016; Nielsen etĀ al., 2017). However, while the absorption of low ions becomes broader and stronger towards the galaxy minor axes, a result attributed to gas kinematically disturbed by outflows (Bordoloi etĀ al., 2011; Martin etĀ al., 2019; Schroetter etĀ al., 2019), the OĀ VI velocity spread does not vary with the sightline azimuthal angle and galaxy disk inclination (Nielsen etĀ al., 2017). Furthermore, for sightlines near the galaxy major axes, the cool gas (traced by MgĀ II) corotates with the galaxy disk and often at sub-centrifugal velocities (Steidel etĀ al., 2002; Kacprzak etĀ al., 2010, 2011; Ho etĀ al., 2017; Zabl etĀ al., 2019), suggesting that the gas is infalling (Ho & Martin, 2020). In apparent conflict with these results, however, OĀ VI studies have not shown such a correlation (Kacprzak etĀ al., 2019, 2025).
Given the spatially extended distribution of O VI compared to the low ions, it is perhaps unsurprising that the O VI kinematics do not seem to correlate with disk rotation as the low-ionization-state counterparts. In particular, because the CGM angular momentum remains roughly constant with radius, the outer CGM is expected to have a lower mean rotation velocity, resulting in a weaker correlation with the disk rotation. To date, however, very few observational studies have focused on the kinematics comparison between O VI and the galaxy disk (Kacprzak et al., 2019; Nateghi et al., 2024; Kacprzak et al., 2025), which is in contrast to the large number of works in previous decades that observationally confirmed the corotation between the low-ionization-state gas and the disk (Steidel et al., 2002; Kacprzak et al., 2010, 2011; Bouché et al., 2013, 2016; Keeney et al., 2013; Chen et al., 2014; Diamond-Stanic et al., 2016; Ho et al., 2017; Martin et al., 2019; Zabl et al., 2019; Ho & Martin, 2020; Lopez et al., 2020; Tejos et al., 2021).
In this paper, we build a sample of star-forming galaxies with quasar sightlines near the galaxy major axes and exclude galaxies with face-on disks. With quasar spectroscopy from the Cosmic Origin Spectrograph (COS) on the Hubble Space Telescope (HST) and galaxy rotation curve measurements, we analyze and compare the kinematics of the CGM traced by OĀ VI and low ions (e.g., SiĀ II, SiĀ III) with the galaxy disk rotation. We describe the sample selection, data reduction, and measurements in SectionĀ 2. In SectionĀ 3, we present the results of the kinematics comparison among OĀ VI gas, low-ionization-state gas, and disk rotation. We interpret our results and discuss their implications in SectionĀ 4. Finally, SectionĀ 5 summarizes the conclusions. Throughout the paper, we adopt the cosmology from Planck Collaboration etĀ al. (2016), with , , , and .
@@ 2 Data and Observations
We present the analyses of 18 galaxyāquasar pairs to study the kinematics of OĀ VI and low-ionization-state absorbers and their relationship with galaxy disk rotation. We conducted quasar spectroscopy with HST/COS, including new spectra from our HST program and archival observations. We also acquired galaxy spectra from multiple ground-based telescopes. We describe our sample selection in SectionĀ 2.1. SectionĀ 2.2 discusses the galaxy spectroscopic observations and rotation curve measurements, and SectionĀ 2.3 describes the galaxy properties. In SectionĀ 2.4, we describe the HST/COS observations and measurements of absorption-line systems.
@@ 2.1 Sample of GalaxyāQuasar Pairs
Our study focuses on star-forming galaxies with quasar sightlines near the major axes of the galaxies. The galaxyāquasar pairs were selected from published CGM studies with known galaxy disk orientation; the disk geometry was obtained either from published work or from the PhotoObj table of the Sloan Digital Sky Survey (SDSS; Blanton etĀ al., 2017). Then, we restricted our sample to galaxies with inclined disks of inclination angles of ā, and each sightline lies within 45ā from the galaxy major axis, i.e., with azimuthal angle ā. This selection with inclined disks and major-axis sightlines excludes regions of the CGM proven to be disturbed by galactic outflows (SectionĀ 1). Furthermore, we avoided sightlines near the galaxy minor axes because of the small projected rotation speed from a rotating disk component, which would make it challenging to distinguish rotation from other potential velocity components (e.g., conical outflows). The same reason also led us to exclude face-on disks, in addition to their large uncertainties in the disk position angles and sightline azimuthal angles.
With our sample selection criteria, we selected our target galaxies from 120 unique, galaxyāquasar pairs from the COS-Halos survey (Tumlinson etĀ al., 2013; Werk etĀ al., 2013), the COS-Burst survey (Heckman etĀ al., 2017), the Multiphase Galaxy Halos survey (Kacprzak etĀ al., 2019), and the Quasars Probing Galaxies (QPG) program (Martin etĀ al., 2019), among which 85% of the galaxy candidates were star-forming. Except for the 50 star-forming galaxies in the QPG program, the other three CGM studies were conducted with HST/COS. Because our work focuses on the kinematics of the OĀ VI gas instead of the detection rate, we required OĀ VI being detected around the 52 star-forming galaxies from those three CGM surveys, reducing our OĀ VI-detected galaxy candidates to 32. Then, we imposed our disk inclination and azimuthal angle cuts, and also visually inspected the images (from the survey papers and/or the SDSS -band images) to ensure the galaxies have reasonable disk geometry measurements. We ended up with 11 galaxies from the three surveys, four of which had no rotation curve data at the time when the galaxies were selected (SectionĀ 2.2.1). As for the QPG program, it focused on the MgĀ II kinematics around star-forming disk galaxies with disk geometry and rotation curve measurements readily available (Ho etĀ al., 2017; Martin etĀ al., 2019). We selected galaxyāquasar pairs that satistied our criteria on azimuthal angle and disk inclination, regardless of whether or not MgĀ II was detected. Then, we selected seven quasars bright enough in the far-UV (FUV) with for new HST/COS observations (SectionĀ 2.4).
The final compiled sample consists of 18 galaxyāquasar pairs with a median galaxy redshift of . TablesĀ 1 and 2 specify the coordinates and properties of galaxies, respectively, and TableĀ 3 lists the coordinates of the paired quasars. FigureĀ 1 shows the location of the quasar sightlines relative to the projected major and minor axes of the galaxies on the sky. The impact parameter of the quasar sightlines span between 20 and 280 kpc. The majority of the sightlines intersect the galaxies within 100 kpc, which corresponds to less than half of the virial radius () as indicated by the colors of the markers (see SectionĀ 2.3).
@@ 2.2.1 MMT Blue and Red Channel Observations
We conducted new galaxy spectroscopic observations using the MMT Blue and Red Channel Spectrographs. On February 28, 2020, we observed the galaxies with the Blue Channel and configured it with the 832 lines mm-1 grating blazed at 7790 Ć . On January 9, 2021, we used the Red Channel with the 1200 lines mm-1 grating blazed at 7700 Ć . For both observations, we applied the blue blocking filter R-63 to block the higher order light. We used the default detector binning, which provided a pixel scale of 03 for both channels and a wavelength dispersion of 0.72 Ć pixel-1 and 0.80 Ć pixel-1 for the Blue and Red Channel, respectively. We observed the target galaxies by aligning the 10-wide long-slit with the galaxy major axes. For both channels, the 10-wide slitlet provided a spectral resolution of 75ā85 km s-1 full width at half maximum intensity (FWHM).
@@
From the two-dimensional galaxy spectra, we measured the galaxy rotation curves from the emission lines (SectionĀ 2.2.4). We also extracted an one-dimensional spectrum for each galaxy and determined the galaxy systemic redshift from the redshifted wavelengths of the , [NĀ II] 6548, 6583, and [SĀ II] 6716, 6731 emission lines.
@@ 2.2.2 Keck Observations
The Keck/LRIS observations were described in detail in Ho etĀ al. (2017) and Martin etĀ al. (2019). The LRIS double spectrograph was configured using the D500 dichroic, 1200 lines mm-1 grism blazed at 3400Ć (LRISb) and the 900 lines mm-1 grism blazed at 5500Ć (LRISr). The galaxy emission lines for rotation curve measurements were captured by LRISr. The detectors were binned by two in both spatial and dispersion directions, which produced a spatial scale of and a dispersion of 1.06 Ć pixel-1. The 10-wide long-slit was either aligned with the position angle of the major axis of the target galaxies, or masks of 10-wide slitlets were designed to include nearby galaxies in the field. The 10-wide slitlets provided an average resolution of 75ā105 km s-1 FWHM. The data reduction procedures for the LRISr spectra were explained in Ho etĀ al. (2017) and were similar to those for reducing the MMT spectra described in SectionĀ 2.2.1.
The Keck/ESI observations were described in detail in Kacprzak etĀ al. (2019). In brief, the galaxies were observed using the 10-wide, 20ā²ā²-long ESI slit, and the slitlet was aligned with the major axes of individual galaxies. The detector was binned by two in both the spatial and spectral directions. This produced a dispersion of 22 km s-1 pixel-1 and pixel scales of 027 to 034 over the echelle orders that covered the galaxy optical emission lines. The resulting spectral resolution from the 10-wide slitlet was 90 km s-1 FWHM. Following the data reduction procedures as outlined in Kacprzak etĀ al. (2019), we re-reduced the ESI spectra and re-examined the galaxy emission lines. The new two-dimensional spectra were adopted to measure the galaxy rotation curves (SectionĀ 2.2.4).
@@ 2.2.3 APO DIS Observations
For galaxies J142459+38211 and J155505+362848, Ho etĀ al. (2017) and Martin etĀ al. (2019) conducted long-slit spectroscopy with APO/DIS to measure the galaxy rotation curves from the emission line. The 15-wide long-slit was aligned with the major axis of each galaxy. The blue and red channels of DIS were configured with the 1200 lines mm-1 blazed at 4400Ć and 7300Ć , respectively. The emission was detected by the red spectra, which had a spatial scale and dispersion of 040 and 0.58 Ć pixel-1, respectively. The DIS spectra were reduced as discussed in Ho etĀ al. (2017) and followed the same procedures as we reduced the MMT spectra. The spectral resolution was measured using the arclampsā lines as 50 km s-1 FWHM.
@@ 2.2.4 Galaxy Rotation Curves
For each galaxy, we project the models onto the slit, convolve these projected models with a Gaussian model for the seeing disk, and compare them to the PV maps measured from the galaxy emission lines. We minimize the fit residual to estimate and for individual galaxies. In the middle column of FigureĀ 2, the seeing-convolved model for each galaxy before and after the projection are shown using the black dashed and cyan solid curves, respectively. TableĀ 2 lists the adopted and for individual galaxies.